# Leveraging Anthropometric Measurements to Improve Human Mesh Estimation and Ensure Consistent Body Shapes

### v1

@@ Appendix C Evaluating A2B Models
We measure two types of errors to evaluate the performance of our A2B models. The first type ($\beta$ error) shows the error if we take the GT $\beta$ parameters, derive anthropometric measurements (B2A), input them into the A2B models and evaluate the MSE of the predicted $\beta$ parameters. The second type ($A$ error) calculates B2A from the predicted $\beta$ parameters and evaluates the mean difference between the GT and predicted anthropometric measurements (all 36) in mm. These evaluations are a kind of cycle consistency evaluation for A2B and B2A. Figure 8 provides a visualization of the evaluation scheme. The part that is also included in the training is highlighted.
@@
Figure 8: Visualization of the A2B evaluation and training procedures. The training part is highlighted with thicker arrows.
Appendix D Keypoint Selection for fit3D
@@
We use the fit3D (Fieraru et al., 2021) dataset for our evaluations, since this is the only sports dataset with public SMPL-X annotations. We evaluate on the SMPL-X joints, since these are trivial to obtain from SMPL-X meshes and there is no regressor available for the fit3D annotated 3D joints. SMPL-X has 144 defined joints. Since our focus is mainly on the body and not on the hands and face, we remove most of these joints. In the end, we select a subset of 37 SMPL-X joints: pelvis, left hip, right hip, spine1, left knee, right knee, spine2, left ankle, right ankle, spine3, left foot, right foot, neck, left collar, right collar, head, left shoulder, right shoulder, left elbow, right elbow, left wrist, right wrist, left index, left thumb, right index, right thumb, left big toe, left small toe, left heel, right big toe, right small toe, right heel, right eye, left eye, right ear, left ear, nose.
Appendix E Generation of Pseudo GT Anthropometric Measurements
@@
As we do not have access to the athletes of ASPset and fit3d to obtain real anthropometric measurements, we need an alternative to kind-of simulate this process. For ASPset, as a first step, we run IK on the GT 3D joint locations. We obtain the necessary anthropometric parameters from the generated meshes with B2A. Then, we use the median values of these measurements as the GT anthropometric values. We call these parameters pseudo GT throughout this paper, since this is not directly the GT, but obtained from IK executed on the GT 3D joint locations and the B2A computation from the created meshes. These parameters are used in this paper to generate the pseudo GT $\beta$ parameters by A2B prediction.
@@
We do not have access to the athletes of the fit3D dataset either. Therefore, we need some kind of GT data to mimic measurements. Obviously, there is no GT available for the official test set evaluation on the evaluation server. We therefore split the official training dataset into a training, validation, and test set for our evaluations, details can be found in the main paper. With this selection, we have GT shape parameters available. We do not use these directly, but apply B2A and use the median measurements over time in order to mimic the measuring process and obtain a single set of anthropometric measurements per person (which is not the case for the provided GT parameters, see Section 3 in the main paper. In real applications, this step is omitted because the anthropometric parameters can be measured directly from the athletes.
Appendix F Finetuning HME Models with Pseudo GT Meshes
@@
Finetuning existing HME models on pure 3D joints datasets is not possible, since they need mesh annotations for training. However, with IK, we can generate pseudo GT meshes. We exemplary test a finetuning of SMPLer-X on ASPset with this approach. Using their finetuning script with 1.6M iterations leads to worse results than the results without finetuning. Therefore, we reduce the number of iterations with early stopping and achieve better results with finetuning only for 32K iterations.
@@
Table 10: MPJPE results in mm for different methods and A2B measurements/median $\beta$ parameters on the test split of ASPset. SMPLer-X FT stands for the best finetuned variant of SMPLer-X (finetuned with the meshes obtained from IK executed on the GT 3D joints). The orig column contains the results without replaced $\beta$ parameters.
model orig. measurements NN m SVR m NN n SVR n median
SMPLer-X 86.02 SMPLer-X 85.89 85.69 86.03 85.99 86.04
SMPLer-X FT 79.09 SMPLer-X FT 78.92 78.88 79.44 79.37 79.44
SMPLer-X FT - GT 65.63 65.84 64.71 64.76 64.79
SMPLer-X FT - SMPLer-X 73.41 73.29 73.65 73.63 73.66
UU IK 67.54 UU 66.92 66.60 67.25 67.12 67.16
UU IK - SMPLer-X 63.80 63.64 63.81 63.78 63.82
UU IK - SMPLer-X FT 69.46 69.27 69.70 69.63 69.69
UU IK - GT 56.44 56.56 55.14 55.19 55.21
We provide a comprehensive summary and visualization of all results on the ASPset dataset in Section G. This includes results of existing HME models, results of our approach, and the finetuning results.
Appendix G Summary of the Results
We execute a multitude of experiments with different combinations of pose and shape parameters. Figure 10 summarizes the results with their pose and shape origins for ASPset. In general, the poses estimated by IK based on the UU results (red branch in Fig. 10) are more precise than the poses estimated by SMPLer-X (light blue branch in Fig. 10). Further, the body shape parameters from our A2B models with GT anthropometric measurements (green boxes in Fig. 10) achieve the best results for all poses. We provide more qualitative examples comparing SMPler-X with this approach in Figure 9. Without access to the GT, all models benefit slightly from A2B model results with the median anthropometric measurements from B2A of the estimated meshes by the respective model (boxes with same color for all three branches in Fig. 10). Moreover, SMPLer-X A2B body shape parameters perform best when analyzing body shapes without GT access (light blue boxes in Fig. 10). Finetuning SMPLer-X with IK created meshes (dark blue branch in Fig. 10) improves the performance of SMPLer-X, although the quality of the body shape deteriorates. This can be seen as by comparing the shapes from SMPLer-X and finetuned SMPLer-X (dark blue and light blue boxes in Fig. 10) with finetuned and IK poses.
Since fit3D is a larger dataset, finetuning UU works better, which further leads to better IK meshes with an MPJPE of 37.02 mm.
Figure 9: Qualitative results of SMPLer-X and our approach for ASPSet example frames.
@@ Appendix G Summary of the Results
Enforcing consistent meshes with GT or IK A2B shape parameters decreases the performance slightly in this case. However, A2B shape parameters achieve slightly better scores than median values. This also holds for OSX and Multi-HMR. Overall, the approach with UU, IK, and A2B body shape parameters achieves an over 33 mm lower MPJPE than any HME model.

### v3

@@ 9 Evaluating A2B Models
We measure two types of errors to evaluate the performance of our A2B models. The first type ($\beta$ error) shows the error if we take the GT $\beta$ parameters, derive anthropometric measurements (B2A), input them into the A2B models and evaluate the MSE of the predicted $\beta$ parameters. The second type ($A$ error) calculates B2A from the predicted $\beta$ parameters and evaluates the mean difference between the GT and predicted anthropometric measurements (all 36) in mm. These evaluations are a kind of cycle consistency evaluation for A2B and B2A. Figure 9 provides a visualization of the evaluation scheme. The part that is also included in the training is highlighted with thicker arrows. The anthropometric error is only used during evaluation.
@@
Figure 9: Visualization of the A2B evaluation and training procedures. The training part is highlighted with thicker arrows. During training, the $\beta$ parameter error is used. For evaluatations, the $\beta$ parameter error and the anthropometric error are calculated.
10 Keypoint Selection for fit3D
@@
We use the fit3D [12] dataset for our evaluations, since this is the only sports dataset with public SMPL-X annotations. We evaluate on the SMPL-X joints, since these are trivial to obtain from SMPL-X meshes and there is no regressor available for the fit3D annotated 3D joints. SMPL-X has 144 defined joints. Since our focus is mainly on the body and not on the hands and face, we remove most of these joints. In the end, we select a subset of 37 SMPL-X joints: pelvis, left hip, right hip, spine1, left knee, right knee, spine2, left ankle, right ankle, spine3, left foot, right foot, neck, left collar, right collar, head, left shoulder, right shoulder, left elbow, right elbow, left wrist, right wrist, left index, left thumb, right index, right thumb, left big toe, left small toe, left heel, right big toe, right small toe, right heel, right eye, left eye, right ear, left ear, nose.
11 Generation of Pseudo GT Anthropometric Measurements
@@
As we do not have access to the athletes of ASPset and fit3d to obtain real anthropometric measurements, we need an alternative to simulate this process. For ASPset, as a first step, we run IK on the GT 3D joint locations. From the generated meshes, we obtain the necessary anthropometric parameters with B2A. Then, we use the median values of these measurements as the GT anthropometric values. We call these parameters pseudo GT throughout this paper, since this is not directly the GT, but obtained from IK executed on the GT 3D joint locations and the B2A computation from the created meshes. These parameters are used in this paper to generate the pseudo GT $\beta$ parameters by A2B prediction.
pose orig. measure NN m SVR m NN n SVR n median
SMPLer-X 86.02 SMPLer-X 85.89 85.69 86.03 85.99 86.04
SMPLer-X FT 79.09 SMPLer-X FT 78.92 78.88 79.44 79.37 79.44
SMPLer-X FT - GT 65.63 65.84 64.71 64.76 -
SMPLer-X FT - SMPLer-X 73.41 73.29 73.65 73.63 73.66
IK-UU 67.54 IK-UU 66.92 66.60 67.25 67.12 67.16
IK-UU - SMPLer-X 63.80 63.64 63.81 63.78 63.82
IK-UU - SMPLer-X FT 69.46 69.27 69.70 69.63 69.69
IK-UU - GT 56.44 56.56 55.14 55.19 -
Table 10: MPJPE results in mm for the test split of ASPset. Results are given for different methods and replaced $beta$ parameters with A2B results (columns NN/SVR) or the median of the original $\beta$ parameters from the model noted in the measure column. SMPLer-X FT stands for the best finetuned variant of SMPLer-X (finetuned with the meshes obtained from IK executed on the GT 3D joints). The orig column contains the results without replaced $\beta$ parameters. We highlight the best result for each model and the best option for the combination of IK-UU pose and SMPLer-X $\beta$ parameters, since this combination outperforms the original IK-UU result, too.
@@
We do not have access to the athletes of the fit3D dataset either. Therefore, we need some kind of GT data to mimic measurements. Obviously, there is no GT available for the official test set evaluation on the evaluation server. We therefore split the official training dataset into a training, validation, and test set for our evaluations. We perform a leave-one-out cross validation, therefore all eight athletes from the official training dataset are used in our evaluation. With this selection, we have real GT shape parameters available. We do not use these directly, since this would skip the measuring process that is needed in real applications. Further, the GT data is not consistent (see Section 3 in the main paper). Therefore, we apply B2A and use the median measurements over time in order to simulate the measuring process and obtain a single set of anthropometric measurements per person. In real applications, this step is omitted because the anthropometric parameters can be measured directly from the athletes before starting the recording.
We consider this strategy as a valid method for evaluations, since our main goal is to improve the HME performace as much as possible with only marginal overhead. Our main focus is sports, which contains extreme poses that lets existing HME models fail, sometimes even to detect a human at all. Examples can be found in the supplementary videos. As professional athletes are measured anyway, the additional effort for the measurements is negligible in this context.
12 Inverse Kinematics
We use the inverse kinematics approach with a VPoser extension, as proposed in the code by [27], to fit SMPLX meshes to given 3D keypoints. VPoser is a learned prior for human poses, since the raw SMPL-X model definition allows impossible poses for humans. VPoser learned plausible poses from the large AMASS [21] dataset and helps IK to generate only plausible poses. IK learns the best SMPL-X parameters ($\beta$ and $\theta$) that fit the mesh to the given 3D joint locations by minimizing the error between the given joint locations and the regressed joint locations from the mesh. IK is an iterative algorithm and adjusts the pose and the shape parameters with a gradient descent minimization approach in each step. Besides the already described joint error, IK further penalizes abnormal poses with a VPoser error and extreme body shapes with a $\beta$ parameter error. Therefore, the total loss for IK can be described as:
[…]
We execute IK per frame, which results in a slight jitter in between the frames, but leads to more accurate joint positions. Since IK needs multiple iterations to adjust the standard T-pose parameters to achieve a pose that is roughly close to the desired UU pose, we speed up the process by initializing the pose and shape parameters with the result from the previous frame if available. This also enhances the final result slightly. We acknowledge that IK is relatively slow regarding the runtime, but our main focus is the precision. For sport analysis, which is our focus, the runtime is not critical, but a very precise result is crucial.
13 Finetuning HME Models with Pseudo GT Meshes
@@
Finetuning existing HME models on pure 3D joints datasets is not possible, since they need mesh annotations for training. However, with IK, we can generate pseudo GT meshes. We exemplary test a finetuning of SMPLer-X on ASPset with this approach. Experiments show that using their finetuning script with 1.6M iterations leads to worse results than the results without finetuning. Therefore, we reduce the number of iterations with early stopping and achieve better results with finetuning only for 32K iterations.
@@
We provide a comprehensive summary and visualization of all results on the ASPset dataset in Section 14. This includes results of existing HME models, results of our approach, and the finetuning results.
14 Summary of the Results
We execute a multitude of experiments with different combinations of pose and shape parameters. Figure 10 summarizes the results with their pose and shape origins for ASPset. In general, the poses estimated by IK based on the UU results (red branch in Fig. 10) are more precise than the poses estimated by SMPLer-X (light blue branch in Fig. 10). Further, the body shape parameters from our A2B models with GT anthropometric measurements (green boxes in Fig. 10) achieve the best results for all poses. We provide more qualitative examples comparing SMPler-X with this approach in the supplementary video. Without access to the GT, all models benefit slightly from A2B model results with the median anthropometric measurements from B2A of the estimated meshes by the respective model (boxes with same color for all three branches in Fig. 10). Moreover, SMPLer-X A2B body shape parameters perform best when analyzing body shapes without GT access (light blue boxes in Fig. 10). Finetuning SMPLer-X with IK created meshes (dark blue branch in Fig. 10) improves the performance of SMPLer-X, although the quality of the body shape deteriorates. This can be seen as by comparing the shapes from SMPLer-X and finetuned SMPLer-X (dark blue and light blue boxes in Fig. 10) with finetuned and IK poses.
Since fit3D is a larger dataset, finetuning UU works better, which further leads to better IK meshes with an MPJPE of 37.02 mm. Enforcing consistent meshes with GT or IK A2B shape parameters decreases the performance slightly in this case. However, A2B shape parameters achieve slightly better scores than median values. This also holds for OSX and Multi-HMR. Overall, the approach with UU, IK, and A2B body shape parameters achieves an over 33 mm lower MPJPE than any HME model. The same also holds for the MVE, which can be improved by over 30 mm with our approach. The scores can be found in the main paper.
We provide two videos in the supplementary material that show qualitative results for ASPset and fit3D. Figure 11 shows one example visualization for both datasets. We include the GT and predicted meshes in the fit3d visualization and display the GT and estimated body shapes in T-pose right next to each other. For ASPset, we visualize the estimated meshes and the GT and estimated joints, since we do not have GT meshes here.
@@ 14 Summary of the Results
Figure 11: Example frames from our supplementary videos. It shows qualitative results of our approach compared to MultiHMR for fit3d (left) and qualitative results of SMPLer-X and our approach for example frames from ASPSet (right). In fit3d visualizations, we display the GT meshes in green and the estimated meshes in grey. The GT joints are also displayed in green while the estimated joints from our model are visualized in blue blue. The MultiHMR joints are shown in red. Corresponding joints are connected. We display the exact MPJPE values in the top left of each frame. Recall that the visualization is in 2D, but the evaluation is in 3D. Therefore, sometimes the MPJPE values may seem odd. In the lower part, we show the estimated body shapes in T-pose. The GT body shape is shown in green and the estimated body shape from our model in blue. The MultiHMR body shape is shown in red. For ASPset visualizations, we display the estimated meshes and the GT and estimated joints. GT joints are shown in green, estimated joints from our model in blue. and the SMPLer-X joints in red. Corresponding joints are connected. In the lower part, we show the GT and estimated joints in the same way, but without the mesh and image to reduce distraction. We further display the MPJPE values.